COMPUTATION OF INFRARED SPECTROSCOPIC FEATURES USING SYMMETRY

\underline{J. VÁZQUEZ}, J. F. STANTON, Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, University of Texas at Austin, Austin, TX 78712.

An efficient algorithm for the prediction of anharmonic infrared features of symmetric top molecules in the framework of second-order vibrational perturbation theory is presented. Higher-order derivatives of the energies are obtained by numerical differentiation of analytic second derivatives in conjuction with symmetry projection which leads to a major reduction of the computational demand. The performance of the implementation is illustrated by several exemples.